Question 2 options:

Given the value for the universal gravitational constant, the average value for the acceleration due to gravity at the surface of Earth, and the radius of Earth, Earth’s mass is ____x1024kg. 



1
Expert's answer
2021-11-22T16:20:33-0500

"F=GMm/r^2\to g=GM/r^2\to"


"M=gr^2/G=9.81\cdot6371000^2/(6.67\cdot10^{-11})=5.97\cdot10^{24}\ (kg)" . Answer